OpenStudy (anonymous):

Use the formula for computing future value using compound interest to determine the value of an account at the end of 6 years if a principal amount of $2,500 is deposited in an account at an annual interest rate of 7% and the interest is compounded daily. (Assume there are 365 days in a year) The amount after 6 years will be $?

OpenStudy (anonymous):

is it 383,250?

OpenStudy (anonymous):

385,750 ?

OpenStudy (anonymous):

I don't know the formula for this.. If you tell me the formula, then I might help you...

OpenStudy (anonymous):

I have a bunch of formulas... idk what one to use? there is.. I=prt A=P(1+rt) P+I/n A=P(1+(r/m))^n

OpenStudy (anonymous):

I think we should go with last one...

OpenStudy (anonymous):

How do I use it.. I have the rate.. which is 0.07 Principal is 2500... what is m and n?

OpenStudy (anonymous):

A=2500(1+(0.07)(6))? or is time 365*6= 2190?

OpenStudy (anonymous):

A=2500(1+(0.07)(2190)

OpenStudy (anonymous):

No, we will go with last one....

OpenStudy (anonymous):

last formula or the 2190 one?

OpenStudy (anonymous):

Last formula... Wait, I explain the formula first to you... Please wait...

OpenStudy (anonymous):

okay.

OpenStudy (anonymous):

Formula is: \[A = P(1 + \frac{r}{n})^{n.t}\] This is the formula for calculating amount in case of Compound Interest...

OpenStudy (anonymous):

Here, A = Amount (that we have to calculate) P = Principle Amount.. (P = 2500) r = Annual Rate of Interest (in Decimal) (r = 0.07) n = Number of Times the Interest is compounded per year..(n = 365) t = Amount of Money accumulated after n years, Including Interest.. (t = 6) Put these values in the formula given above and solve for A.. Can you do it??

OpenStudy (anonymous):

3804.7507

OpenStudy (anonymous):

so to the nearest cent would that be 3804.75

OpenStudy (anonymous):

How did you solve it show me the steps so that I can verify it...

OpenStudy (anonymous):

A=2500(1+(0.07/365)^365*6 A=2500(1+(0.07/365)^2190 A=2500(1.5219) A=3804.75

OpenStudy (anonymous):

Thanks for solving it for me.. See, I also don't know the formula but now I know.. Thanks for increasing my knowledge...

OpenStudy (anonymous):

lol, so do you think I did it right?

OpenStudy (anonymous):

Yes I think so...

OpenStudy (anonymous):

okay thanks.